# Optimizing Breast Cancer Detection in Mammograms: A Comprehensive Study of Transfer Learning, Resolution Reduction, and Multi-View Classification

### v1

@@ 4.2.2 Learn-to-Resize Classifier (LRC)
DenseNet121 224 0.6712$\pm$0.0208 0.6518$\pm$0.0222
ResNeXt-50-32x4d 224 0.6735$\pm$0.0042 0.6682$\pm$0.0220
ConvNeXt-Small 224 0.6734$\pm$0.0080 0.6709$\pm$0.0219
DenseNet169 224 0.6718$\pm$0.0178 0.6850$\pm$0.0216
DenseNet201 224 0.6848$\pm$0.0018 0.6855$\pm$0.0216
ResNet-18 224 0.6847$\pm$0.0030 0.6884$\pm$0.0216
ConvNeXt-Base 224 0.6791$\pm$0.0076 0.6892$\pm$0.0216
EfficientNet-B4 384 0.7501$\pm$0.0274 0.7113$\pm$0.0211
EfficientNet-B0 224 0.7494$\pm$0.0228 0.7174$\pm$0.0209
ResNet-50 224 0.7188$\pm$0.0082 0.7272$\pm$0.0207
MobileNetV2 224 0.7096$\pm$0.0269 0.7297$\pm$0.0206
ResNet-101 224 0.7230$\pm$0.0226 0.7532$\pm$0.0200
EfficientNet-B1 240 0.7477$\pm$0.0063 0.7547$\pm$0.0200
EfficientNet-B3 300 0.7613$\pm$0.0094 0.7564$\pm$0.0199
ConvNeXt-Tiny 224 0.7200$\pm$0.0273 0.7574$\pm$0.0199
MNASNet-1.0 224 0.6979$\pm$0.0464 0.7626$\pm$0.0197
EfficientNet-B2 260 0.7497$\pm$0.0345 0.7690$\pm$0.0195
ConvNeXt-Base (22k) 224 0.7399$\pm$0.0393 0.7707$\pm$0.0195
EfficientNetV2-M 384 0.7335$\pm$0.0297 0.7749$\pm$0.0193
EfficientNetV2-S (21k) 300 0.7719$\pm$0.0122 0.7779$\pm$0.0192
EfficientNetV2-S 300 0.7475$\pm$0.0342 0.7890$\pm$0.0188
MobileNetV3-Large 224 0.7724$\pm$0.0194 0.7958$\pm$0.0186
@@ 4.2.3 Conclusions on Resizing
Hypothesis test described on page 4.1.5 reveals that patch-based classification (PBC) is significantly superior to LRC, with a p-value of 0.0202. However, there is limited statistical evidence to conclude that PBC outperforms FRC, as the p-value for this comparison is 0.1828.
@@
4.3 Two-View Classifier
In this section, we aim to address the question (4) listed in Section 3.
@@ 4.3.1 Architecture of Two-View Classifier
We assembled the network as in Fig. 2 making another transfer learning with the weights obtained from the best single-view models of section 4.1. We obtained the AUCs described in Table 9.
@@
The best two-view classifier was achieved using the direct classification (DC) approach with the EfficientNet-B3 base model, yielding an AUC of 0.8643. This represents the highest AUC recorded for classifying two-view exams in the CBIS-DDSM dataset under its original training/testing split. This result surpasses the AUC of 0.8418 reported in [7] using EfficientNet-B0, with some statistical evidence of superiority (p-value = 0.0821, using hypothesis test described on page 4.1.5). We attribute this improvement to the use of a base model trained with more advanced and modern techniques, as discussed in Section 4.1.5. Using Test Time Augmentation (TTA), we achieved an AUC of 0.8658, compared to the AUC of 0.8483 reported in [7]. The hypothesis test yielded a p-value of 0.1366, indicating weak statistical evidence that our result is superior to the previous.
@@ 4.3.2 Average or Maximum Operation
EfficientNet-B3 (DC) 0.8643$\pm$0.0241 0.8420$\pm$0.0257 0.8426$\pm$0.0257
We conducted the DeLong test [^4] An AUC comparison method developed by DeLong et al. [20]. We used the one-tailed fast version proposed by Sun, Xu, and Xu [21], with the implementation available at https://github.com/yandexdataschool/roc_comparison, computing the unadjusted AUC covariance. to compare the two-view model with the mean and maximum operations (Table 10). The resulting p-values were 0.0280 and 0.0286, respectively.
@@
5 Architecture exploration in VinDr-Mammo
In this section, we will repeat the experiments using the VinDr-Mammo dataset, which consists entirely of Full-Field Digital Mammographies (FFDMs), to answer the question (5) listed in Section 3. Unlike CBIS-DDSM, this dataset does not include biopsy-confirmed benign or malignant labels. Instead, it provides Bi-RADS (Breast Imaging Reporting and Data System) and other annotations, uniformly distributed across 4,000 training exams and 1,000 test exams.
@@ 5 Architecture exploration in VinDr-Mammo
5.1 Single View Classifier
@@ 5.1.1 Patch Classifier and Base-Model
To answer questions (1) and (2) of Section 3, we prepared the patches similar to Section 4.1.2, considering four categories: background and the lesions with Bi-Rads 3, 4 and 5. The backbone model with the best performance was ConvNeXt_Base_22k (Table 13).
@@ 5.1.5 Conclusions on Base Models
5.2 Classifiers with Resized Images
To answer the question (3) listed in Section 3 for VinDr-Mammo, we downsample mammograms to half their size before classifying them.
@@ 5.2.3 Conclusions on Resizing
5.3 Two-View Classifier
To address question (4), we followed the same network creation and training process outlined in Section 4.3. We constructed a two-view classifier using the ConvNeXt_Base_22k network, initializing it with weights from the single-view PBC approach. Additionally, we processed each view (CC and MLO) individually using the best-performing classifier from the previous section (ConvNeXt_Base_22k) and combined the results using both average and maximum operations. The results are shown in Table 18.
@@ 6 Conclusions
Single-view classifiers using PBC and DC approaches yielded very similar AUCs on the CBIS-DDSM dataset. Given that PBC is more complex and computationally intensive compared to DC, we recommend using ImageNet weights directly for classifying low-quality mammograms. On the VinDr-Mammo dataset, the PBC approach achieved a significantly higher AUC compared to DC. Therefore, despite the increased training time, we recommend PBC approach for building classifiers for high-quality mammograms.
We computed the Pearson correlation coefficients between the performance of the PBC and DC approaches and the base models’ accuracies on ImageNet. The PBC approach showed moderately high correlations, suggesting that its performance improves with the use of more modern networks. In contrast, DC exhibited weak correlations, indicating that using modern base models does not provide significant benefits for this approach.
The best classification results for VinDr-Mammo were achieved using a modern network (ConvNeXt), whereas the top results for CBIS-DDSM were obtained with older networks (EfficientNet family). This suggests that the fine details in fully digital mammograms may necessitate more advanced base models for optimal performance.
The downsampled classifiers (576×448) produced lower AUCs compared to classifiers at the original resolution (1152×896) across both datasets and approaches. Therefore, we do not recommend reducing mammogram resolution by 50% prior to classification. In both datasets, the Fixed Resizing Classifier (FRC) achieved higher AUCs than the Learn-to-Resize Classifier (LRC). Given that FRC is also less computationally demanding than LRC, there is no practical advantage to using the “learn to resize” approach.
For both sets, the two-view classifier was clearly superior to calculating the mean or maximum of the outputs of the CC and MLO individual views.

### v2

@@ 3.3.2 Learn-to-Resize Classifier (LRC)
[…]
@@ 3.3.3 Conclusions on Resizing
Hypothesis test described earlier reveals that patch-based classification (PBC) is significantly superior to LRC, with a p-value of 0.0202. However, there is limited statistical evidence to conclude that PBC outperforms FRC, as the p-value for this comparison is 0.1828.
@@
3.4 Two-View Classifiers in CBIS-DDSM
In this section, we aim to address the top question (4).
@@ 3.4.1 Architecture of Two-View Classifier
We assembled the network as in Fig. 3 making another transfer learning with the weights obtained from the best single-view models. We obtained the AUCs described in Table 9.
@@
The best two-view classifier was achieved using the direct classification (DC) approach with the EfficientNet-B3 base model, yielding an AUC of 0.8643. This represents the highest AUC recorded for classifying two-view exams in the CBIS-DDSM dataset under its original training/testing split. This result surpasses the AUC of 0.8418 reported in [6] using EfficientNet-B0, with some statistical evidence of superiority (p-value = 0.0821, using the hypothesis test already mentioned). We attribute this improvement to the use of a base model trained with more advanced and modern techniques, as discussed earlier. Using Test Time Augmentation (TTA), we achieved an AUC of 0.8658, compared to the AUC of 0.8483 reported in [6]. The hypothesis test yielded a p-value of 0.1366, indicating weak statistical evidence that our result is superior to the previous.
@@ 3.4.2 Average or Maximum Operation
[…]
We conducted the DeLong test [^2] An AUC comparison method developed by DeLong et al. [40]. We used the one-tailed fast version proposed by Sun, et al. [41], with an available implementation [42], for computing the unadjusted AUC covariance. to compare the two-view model with the mean and maximum operations (Table 10). The resulting p-values were 0.0280 and 0.0286, respectively.
@@
3.5 Architecture exploration in VinDr-Mammo
In this section, we will repeat the experiments using the VinDr-Mammo dataset [16], which consists entirely of Full-Field Digital Mammographies (FFDMs), to answer the question (5). Unlike CBIS-DDSM, this dataset does not include biopsy-confirmed benign or malignant labels. Instead, it provides Bi-RADS (Breast Imaging Reporting and Data System) and other annotations, uniformly distributed across 4,000 training exams and 1,000 test exams.
@@ 3.5 Architecture exploration in VinDr-Mammo
3.6 Single View Classifiers in VinDr-Mammo
@@ 3.6.1 Patch Classifier and Base-Model
To answer questions (1) and (2) for VinDr-Mammo dataset, we prepared the patches similar to Section 3.2.2, considering four categories: background and the lesions with Bi-Rads 3, 4 and 5. The backbone model with the best performance was ConvNeXt_Base_22k (Table 13).
@@ 3.6.5 Conclusions on Base Models
3.7 Classifiers with Resized Images in VinDr-Mammo
To answer the question (3) for VinDr-Mammo, we downsample mammograms to half their size before classifying them.
@@ 3.7.3 Conclusions on Resizing
3.8 Two-View Classifiers in VinDr-Mammo
To address question (4), we followed the same network creation and training process outlined in previous sectionshae. We constructed a two-view classifier using the ConvNeXt_Base_22k network, initializing it with weights from the single-view PBC approach. Additionally, we processed each view (CC and MLO) individually using the best-performing classifier from the previous section (ConvNeXt_Base_22k) and combined the results using both average and maximum operations. The results are shown in Table 18.
@@ 4 Conclusions
In this study, we systematically evaluated different training approaches for several deep neural network architectures using the film-based CBIS-DDSM dataset (22 architectures from eight network families) and the digital VinDr-Mammo dataset (seven architectures from six families). For low-quality mammograms (CBIS-DDSM), we observed no significant benefit from patch-based pretraining approach; however, it significantly improved classification accuracy on high-quality mammograms (VinDr-Mammo, p = 0.0013).
A positive correlation emerged between a model’s performance on natural-image classification tasks and mammogram classification when patch-based pretraining was applied (CBIS-DDSM: r = 0.6491; VinDr-Mammo: r = 0.7288). Conversely, without patch-based pretraining, this correlation was negligible or negative, indicating that modern backbone architectures, like ConvNext for VinDr-Mammo, are advantageous primarily when combined with patch-based pretraining.
Reducing mammogram resolution consistently diminished model performance, especially for high-quality digital images. Furthermore, employing a two-view classification approach substantially outperformed single-view classifiers across both datasets (CBIS-DDSM: p-values under 0.0286; VinDr-Mammo: p-values under 0.0030). This strongly supports the adoption of multi-view strategies in future mammogram analysis frameworks.